# Euclid Quick Data Release (Q1)

### v1

@@
The first survey data release by the Euclid mission covers approximately in the Euclid Deep Fields to the same depth as the Euclid Wide Survey. This paper showcases, for the first time, the performance of cluster finders on Euclid data and presents examples of validated clusters in the Quick Release 1 (Q1) imaging data. We identify clusters using two algorithms (AMICO and PZWav) implemented in the Euclid cluster-detection pipeline. We explore the internal consistency of detections from the two codes, and cross-match detections with known clusters from other surveys using external multi-wavelength and spectroscopic data sets. This enables assessment of the Euclid photometric redshift accuracy and also of systematics such as mis-centring between the optical cluster centre and centres based on X-ray and/or Sunyaev–Zeldovich observations. We report 426 joint PZWav and AMICO-detected clusters with high signal-to-noise ratios over the full Q1 area in the redshift range . The chosen redshift and signal-to-noise thresholds are motivated by the photometric quality of the early Euclid data. We provide richness estimates for each of the Euclid-detected clusters and show its correlation with various external cluster mass proxies.
Out of the full sample, 77 systems are potentially new to the literature. Overall, the Q1 cluster catalogue demonstrates a successful validation of the workflow ahead of the Euclid Data Release 1, based on the consistency of internal and external properties of Euclid-detected clusters.
@@
The European Space Agency (ESA) Euclid mission is designed to map the geometry of the Universe in unprecedented detail using optical and near-infrared wide-field imaging and spectroscopy; the completed survey will cover deg2 of the extragalactic sky, observing over 1.5 billion galaxies out to a redshift (Euclid Collaboration: Mellier et al. 2024). Furthermore, Euclid is expected to detect and characterise approximately galaxy clusters and groups (Sartoris et al. 2016; Euclid Collaboration: Adam et al. 2019), driving a tremendous leap in cosmological constraints from joint cluster abundance, clustering, and weak lensing mass estimates (e.g., Maturi et al. 2019; Lesci et al. 2020; Abbott et al. 2020; Costanzi et al. 2020; Romanello et al. 2023).
This paper presents a validation of the Euclid galaxy cluster workflow on the Euclid Quick Release 1 (Q1) data set (Euclid Quick Release Q1 2025), constituting the first survey data released by the mission (Euclid Collaboration: Aussel et al. 2025). The data consists of three non-contiguous fields that cover approximately 63 deg2 – one in the northern and two in the southern hemisphere. We detect clusters and characterise candidates with high signal-to-noise ratio (S/N) across the three fields, favoring a conservative approach during this early phase of Euclid data.
Numerous simulation-based evaluations have been conducted on the accuracy of recovering Euclid cluster masses (e.g., Euclid Collaboration: Ingoglia et al. 2024; Euclid Collaboration: Giocoli et al. 2024), as well as multi-wavelength cluster observables (Euclid Collaboration: Ragagnin et al. 2024). Moreover, extensive tests of the cluster workflow have been performed on the Flagship simulation (Euclid Collaboration: Castander et al. 2024) to better understand the Euclid selection function for clusters (Cabanac et al., in preparation). This paper explores the impact of galaxy selection and quality cuts on the cluster workflow’s ability to robustly detect and characterise galaxy clusters on real Euclid data for the first time.
We examine the distribution in S/N, optical richness and redshift for our cluster detections. In addition, cluster redshift estimates are compared to external photometric and spectroscopic data. We utilise the extensive multi-wavelength coverage at near-infrared, optical, millimetre, and X-ray wavelengths of the Q1 fields in the literature to identify counterparts of the Euclid detections. We evaluate centring offset distributions and study the properties of missed objects.
@@ 2.1 The three Euclid Q1 patches
The Q1 survey area comprises the Euclid Deep Field North (EDF-N), the Euclid Deep Field South (EDF-S), and the Euclid Deep Field Fornax (EDF-F), covering a total area of 63 deg2. Based on the galaxy catalogues used in this analysis, the total area for each field is reported in Fig. 1. The EDF-N is an approximately 23 deg2 circular field located at the northern ecliptic pole. The EDF-F is a 12 deg2 circular field including the Chandra Deep Field South (CDF-S), which has numerous ground- and space-based ancillary observations (as described in Sect. 4). Lastly, the EDF-S is a 28 deg2 field with an extended shape that encompasses two adjacent deep-drilling fields of the Vera C. Rubin Legacy Survey of Space and Time (LSST).
@@ 2.2 The galaxy selection
In addition to these criteria, three sets of additional masks are applied: larger circular masks around bright stars to reject areas that are spoiled by image artefacts in the ground-based images; elliptical masks to cover bright galaxies in the foreground; and masks around extended stellar systems. The last two sets of masks are derived from the masking performed by the DR10 Legacy Survey.666https://www.legacysurvey.org/dr10/external
@@ 2.3 Photometric redshifts
In Fig. 2, we compare the median values of the PDZ to the associated spectroscopic redshift. Statistical errors are computed in the redshift range from 0 to 1.5 where most of the spectroscopic data is concentrated. Defining the normalised offset , where is the PDZ median value and is the spectroscopic value, a relatively stable mean scatter of and a residual bias of () can be observed for the EDF-F (EDF-N). We do not make this comparison for the EDF-S due to the aforementioned much lower number of objects with external spectroscopic counterparts; however, we expect a similar behaviour to EDF-F due to the same external band coverage.
To evaluate how well the shape of the PDZ reflects the measured scatter, we compute the average width of the 15th-to-85th percentiles around the median in redshift bins and compare it to the measured scatter. The result is shown in Fig. 3. On average, the evolution of the PDZ width with redshift is in relatively good agreement with the scatter. However, the PDZ widths are systematically narrower than expected across the entire redshift range by about 65% for the EDF-N and 15% for the EDF-F field. The underestimation of photo- uncertainties, particularly in extreme cases with an effective , can lead to undesired spurious detections if not addressed adequately. To mitigate this issue, we tested two different approaches in the detection algorithms as discussed in Sect. 3.1.
@@ 3.1.1 AMICO
We briefly describe the main features of AMICO (Adaptive Matched Identifier of Clustered Objects), with more details available in Bellagamba et al. (2018) and Maturi et al. (2019). AMICO works with catalogues of galaxy angular positions, magnitudes, and full photometric redshift probabilities, avoiding any explicit selection based on the cluster red sequence. The data are modelled as the sum of the desired cluster contribution and a noise component representing contamination from field galaxies. The cluster filter model includes a Schechter luminosity function (Schechter 1976) and a projected NFW (Navarro, Frenk, & White 1996) radial profile, while the field galaxy distribution at a certain redshift is approximated from the full galaxy sample, weighted by their redshift probabilities. The filter is optimal in the sense that it minimises variance to estimate the signal amplitude, , calculated on a three-dimensional grid of positions and redshifts (with a chosen resolution of 0.01 in both angular position, in degrees, and redshift). Clusters are identified as the peaks with the highest S/N, followed by iterative ‘cleaning’ to remove the contributions of detected clusters, thus reducing blending for refining subsequent detections. This process continues to a minimum S/N, directly linked to the amplitude. For this data set, a regularisation term has been introduced to address galaxies whose photometric redshifts exhibit excessively small uncertainties, indicating potential issues with their redshift estimates. Such galaxies can produce an excess of signal in the AMICO amplitude maps, resulting in potential spurious detections. To avoid this issue, the PDZ of galaxies with photometric redshift uncertainties smaller than , is modelled as a Gaussian with and mode located at . Here, is the median value of the input PDZ.
@@ 3.1.2 PZWav
The PZWav detection algorithm applies a difference-of-Gaussians smoothing kernel that is designed to detect over-densities on cluster scales, while minimizing the impact of larger scale structures on the detection. The code takes as input the positions and full photometric redshift probability distribution functions, PDZ, for each galaxy and generates a data cube in position and redshift containing this information. We refer the reader to Euclid Collaboration: Adam et al. (2019), Werner et al. (2023), and Thongkham et al. (2024) for more detail on the PZWav algorithm. Of particular relevance for the current work, the standard deviations for inner and outer Gaussians in the kernel are set at 300 kpc and 2 Mpc. This choice of scales has been found to effectively remove large-scale structure contributions and isolate clusters in both previous surveys (Thongkham et al. 2024) and tests on the Flagship simulations (Cabanac et al. in preparation). The chosen redshift bin size in the search is set to 0.06. The wider redshift binning compared to AMICO, by construction, regularises the contribution of galaxies with an excessively narrow PDZ, which would otherwise degrade the results.
@@ 3.2 Building the combined catalogue
The cluster detection algorithms were applied to each of the three Q1 fields. The most reliable detections, according to S/N, were selected in the redshift range . We note that the upper redshift limit is lower than the one planned for subsequent data releases (i.e., ), because photometry and photometric redshift estimation still have to be refined for these newly released data; the upper redshift limit for this sample is set by the poorly constrained behaviour of the Euclid photometric redshifts at , as shown in Figs. 2 and 3. In addition, as described in Sect. 2.2, the provisional requirement that all objects in the galaxy catalogues be detected in the VIS instrument necessarily restricts us to lower redshifts than what is anticipated for future data releases. While the principal catalogue is limited to , we nevertheless explore objects at in the joint catalogue in Sect. 4.7. This extension is undertaken with the knowledge of limitations in redshift accuracy, allowing us to assess the reliability of the cluster detections and highlight the most promising candidates in this regime.
@@
For this initial Q1 analysis, we ensured a consistent number density of detections between the two algorithms by adjusting the minimum S/N accordingly. Specifically, we used thresholds of S/N= () for AMICO and S/N= () for PZWav, to guarantee detections per square degree for each algorithm in the EDF-F and EDF-S (EDF-N). The different S/N thresholds applied to the southern and northern fields is due to the different photometric redshift quality, as discussed in Sect. 2.3. The absolute difference in values adopted for AMICO and PZWav is due to different definitions of the S/N by each finder. The S/N thresholds for each of the three fields were determined by visually inspecting a large number of detections to ensure an initial cluster sample that excludes clearly spurious detections. In the following analysis, we prioritise a joint sample of clusters in common between PZWav and AMICO to mitigate known issues in the input galaxy photometric redshifts.
Focusing on the joint, high S/N detections for Q1 allows us to perform an initial internal validation of the cluster workflow for the most reliable detections, enabling us to identify any key issues that may impact the catalogue and subsequent cluster science applications.
To match the cluster detections between AMICO and PZWav, a physical distance of 1 Mpc was chosen, from which an angular search aperture was computed based on the mean cluster redshift of each pair.
We imposed an additional constraint on any redshift difference of to minimise the number of erroneous matches to clusters in projection. If multiple matches were found, the highest S/N target was chosen. The first rows of the resulting joint catalog are presented in Table 2. For each entry, we preserved the positions, redshifts, and S/N s derived by each detection algorithm, whereas the presented richness is computed at the position of the PZWav centers.
@@
The left panel of Fig. 6 compares the redshifts estimated by the two algorithms for the matched catalogue. The pink shaded region illustrates the matching interval, and we see that the redshift difference between matched objects is well within the interval, reinforcing the validity of the matching procedure. The right panel of Fig. 6 shows the internal consistency in the centring performance of the two algorithms.
@@
We compute the richnesses, , using only the photo- assignments for this study. We calculate the richness based on the PZWav cluster centre and redshift, but would obtain similar results by using the AMICO centre and redshift, based on the level of positional and redshift agreement shown in the two panels of Fig. 6. The combined richness and redshift distributions are presented in Fig. 8.
@@ 4.1 Searching for Euclid counterparts in external data
We cross-correlated the joint Q1 cluster catalogue with external meta-catalogues777A meta-catalogue is a combination of different source catalogues with cross-identification of clusters and homogenisation of key quantities, notably redshift and a mass proxy. and individual catalogues. We used meta-catalogues and catalogues primarily selected in X-ray – Meta-Catalogue of X-ray Clusters II (MCXC-II) described in Sadibekova et al. 2024, eROSITA (Bulbul et al. 2024; Kluge et al. 2024) – with the Sunyaev–Zeldovich (SZ) effect – Meta-Catalogue of SZ Clusters (MCSZ) described in Tarrio et al. (in preparation) – and in the optical – Dark Energy Survey (DES) Y1 RedMaPPer (Rykoff et al. 2016; Abbott et al. 2020), and the Abell catalogue (Abell 1958; Abell et al. 1989). We also used the joint X-ray and SZ selected catalogue ComPRASS (Tarrío et al. 2019), the LC2 (Sereno 2015), and the Meta-Catalogue of Cluster Dispersions (MCCD) described in Euclid Collaboration: Melin et al. (in preparation) for clusters with high-quality velocity dispersion measurements.
@@
In total, we found 4/21/85 counterparts in the external catalogues for the EDF-N, EDF-F, and EDF-S fields, respectively. In the EDF-F and the EDF-S, the matches are mainly from RedMaPPer and eROSITA. The EDF-N field is not covered by these two catalogues, which explains the significantly lower number of counterparts that are mainly from the MCXC-II. The total number of detections and matches from the external surveys that fall within the Q1 footprint are listed in Table 1. A detailed breakdown of the individual matches to data sets per Q1 field can be found in Appendix B.

### v2

@@
The first survey data release by the Euclid mission covers approximately in the Euclid Deep Fields to the same depth as the Euclid Wide Survey. This paper showcases, for the first time, the performance of cluster finders on Euclid data and presents examples of validated clusters in the Quick Release 1 (Q1) imaging data. We identify clusters using two algorithms (AMICO and PZWav) implemented in the Euclid cluster-detection pipeline. We explore the internal consistency of detections from the two codes, and cross-match detections with known clusters from other surveys using external multi-wavelength and spectroscopic data sets. This enables assessment of the Euclid photometric redshift accuracy and also of systematics such as mis-centring between the optical cluster centre and centres based on X-ray and/or Sunyaev–Zeldovich observations. We report 426 joint PZWav and AMICO-detected clusters with high signal-to-noise ratios over the full Q1 area in the redshift range . The chosen redshift and signal-to-noise thresholds are motivated by the photometric quality of the early Euclid data. We provide richness estimates for each of the Euclid-detected clusters and show its correlation with various external cluster mass proxies. Due to the limited area and evolving data quality, the sample is not intended to serve as a reference for cosmological applications, but to verify and validate the cluster workflow. Out of the full sample, 77 systems are potentially new to the literature. Overall, the Q1 cluster catalogue demonstrates a successful validation of the workflow ahead of the Euclid Data Release 1, based on the consistency of internal and external properties of Euclid-detected clusters.
@@
The European Space Agency (ESA) Euclid mission is designed to map the geometry of the Universe in unprecedented detail using optical and near-infrared wide-field imaging and spectroscopy; the completed survey will cover deg2 of the extragalactic sky, observing over 1.5 billion galaxies out to a redshift (Euclid Collaboration: Mellier et al. 2025). Furthermore, Euclid is expected to detect and characterise approximately galaxy clusters and groups (Sartoris et al. 2016; Euclid Collaboration: Adam et al. 2019), driving a tremendous leap in cosmological constraints from joint cluster abundance, clustering, and weak lensing mass estimates (e.g., Maturi et al. 2019; Lesci et al. 2022; Abbott et al. 2020; Costanzi et al. 2021; Romanello et al. 2025).
This paper presents a validation of the Euclid galaxy cluster workflow on the Euclid Quick Release 1 (Q1) data set (Euclid Quick Release Q1 2025), constituting the first survey data released by the mission (Euclid Collaboration: Aussel et al. 2025). The data consists of three non-contiguous fields that cover approximately 63 deg2 – one in the northern and two in the southern hemisphere. We detect clusters and characterise candidates with high signal-to-noise ratio (S/N) across the three fields, favouring a conservative approach during this early phase of Euclid data.
Numerous simulation-based evaluations have been conducted on the accuracy of recovering Euclid cluster masses (e.g. Euclid Collaboration et al. 2025b; Euclid Collaboration: Giocoli et al. 2024), as well as multi-wavelength cluster observables (Euclid Collaboration: Ragagnin et al. 2025). Moreover, extensive tests of the cluster workflow have been performed on the Flagship simulation (Euclid Collaboration et al. 2025a) to better understand the Euclid selection function for clusters (Cabanac et al., in preparation). This paper explores the impact of galaxy selection and quality cuts on the cluster workflow’s ability to robustly detect and characterise galaxy clusters on real Euclid data for the first time.
We examine the distribution in S/N, optical richness and redshift for our cluster detections. In addition, cluster redshift estimates are compared to external photometric and spectroscopic data. We utilised the extensive multi-wavelength coverage at near-infrared, optical, millimetre, and X-ray wavelengths of the Q1 fields in the literature to identify counterparts of the Euclid detections. We evaluate centring offset distributions and study the properties of missed objects.
@@ 2.1 The three Euclid Q1 patches
The Q1 survey area comprises the Euclid Deep Field North (EDF-N), the Euclid Deep Field South (EDF-S), and the Euclid Deep Field Fornax (EDF-F), covering a total area of 63 deg2. Based on the galaxy catalogues used in this analysis, the total area for each field is reported in Fig. 1. The EDF-N is an approximately 23 deg2 circular field located at the northern ecliptic pole. The EDF-F is a 12 deg2 circular field including the Chandra Deep Field South (CDF-S), which has numerous ground- and space-based ancillary observations (as described in Sect. 4). Lastly, the EDF-S is a 28 deg2 field with an extended shape that encompasses two adjacent deep-drilling fields of the Vera C. Rubin Legacy Survey of Space and Time (LSST). The exact spatial distribution of the fields is illustrated in Figure 1 of the Q1 overview paper, Euclid Collaboration: Aussel et al. (2025).
@@ 2.2 The galaxy selection
A more detailed description of the individual flags used in the galaxy selection can be found in Euclid Collaboration: Romelli et al. (2025). In addition to these criteria, three sets of additional masks are applied: larger circular masks around bright stars to reject areas that are spoiled by image artefacts in the ground-based images; elliptical masks to cover bright galaxies in the foreground; and masks around extended stellar systems. The last two sets of masks are derived from the masking performed by the DR10 Legacy Survey.666https://www.legacysurvey.org/dr10/external
@@ 2.3 Photometric redshifts
In Fig. 2, we compare the median values of the PDZ to the associated spectroscopic redshift. Statistical errors are computed in the redshift range from 0.05 to 1.5 where most of the spectroscopic data is concentrated. Defining the normalised offset , where is the PDZ median value and is the spectroscopic value, a relatively stable mean scatter of and a residual bias of () can be observed for the EDF-F (EDF-N). We do not make this comparison for the EDF-S due to the aforementioned much lower number of objects with external spectroscopic counterparts; however, we expect a similar behaviour to EDF-F due to the same external band coverage.
In this study, PDZ s are also critical to associate errors to the photometric redshift measurements, which, in turn, are used to weight galaxies in the cluster detection algorithms presented in Sect. 3.1. To evaluate how well the shape of the PDZ s reflect the ‘true’ photometric redshift errors, we compute the average width of the 15th-to-85th percentiles around the median and compare it to the average offset of the PDZ median value with its associated spectroscopic redshift, where available. This test is performed in photometric redshift bins from 0.2 to 2 which corresponds to the Euclid nominal range for cluster searches. The result is shown in Fig. 3. On average, the evolution of the PDZ width with the median photometric redshift is in good agreement with the scatter. However, the PDZ widths are systematically narrower than expected across the entire redshift range by about 65% for the EDF-N and 15% for the EDF-F field. The underestimation of photo- uncertainties, particularly in extreme cases with an effective , can lead to undesired spurious detections if not addressed adequately. To mitigate this issue, we tested two different approaches in the detection algorithms, discussed in Sect. 3.1.
@@ 3.1.1 AMICO
We briefly describe the main features of AMICO (Adaptive Matched Identifier of Clustered Objects), with more details available in Bellagamba et al. (2018) and Maturi et al. (2019). AMICO works with catalogues of galaxy angular positions, magnitudes, and full photometric redshift probabilities, avoiding any explicit selection based on the cluster red sequence. The data are modelled as the sum of the desired cluster contribution and a noise component representing contamination from field galaxies. The cluster filter model includes a Schechter luminosity function (Schechter 1976) and a projected NFW (Navarro, Frenk, & White 1996) radial profile, while the field galaxy distribution at a certain redshift is approximated from the full galaxy sample, weighted by their redshift probabilities. The filter is optimal in the sense that it minimises variance to estimate the signal amplitude, , calculated on a three-dimensional grid of positions and redshifts (with a chosen resolution of 0.01 in both angular position, in degrees, and redshift). Clusters are identified as the peaks with the highest S/N, followed by iterative ‘cleaning’ to remove the contributions of detected clusters, thus reducing blending for refining subsequent detections. This process continues to a minimum S/N, directly linked to the amplitude. For this data set, a regularisation term has been introduced to address galaxies whose photometric redshifts exhibit excessively small uncertainties, indicating potential issues with their redshift estimates. Such galaxies can produce an excess of signal in the AMICO amplitude maps, resulting in potential spurious detections. To avoid this issue, the PDZ of galaxies with photometric redshift uncertainties smaller than , is modelled as a Gaussian with and mode located at . Here, is the median value of the input PDZ. We observed an excess in low redshift detections (below ) prior to the regularisation in AMICO, subsequently obtaining comparable numbers to PZWav. Going forwards for DR1, we plan to implement the same criteria, but with a sigma value modelled as a polynomial function fitted to the data—produced from a new release of the photometric redshift algorithm—rather than a scalar value. A similar procedure will be applied to PZWav.
@@ 3.1.2 PZWav
The PZWav detection algorithm applies a difference-of-Gaussians smoothing kernel that is designed to detect over-densities on cluster scales, while minimising the impact of larger scale structures on the detection. The code takes as input the positions and full photometric redshift probability distribution functions, PDZ, for each galaxy and generates a data cube in position and redshift containing this information. We refer the reader to Euclid Collaboration: Adam et al. (2019), Werner et al. (2023), and Thongkham et al. (2024) for more detail on the PZWav algorithm. Of particular relevance for the current work, the standard deviations for inner and outer Gaussians in the kernel are set at 300 kpc and 1.2 Mpc. This choice of scales has been found to effectively remove large-scale structure contributions and isolate clusters in both previous surveys (Thongkham et al. 2024) and tests on the Flagship simulations (Cabanac et al. in preparation). Purity and completeness evaluations of the PZWav algorithm are also performed in these two papers. Moreover, previous cluster detection with PZWav has found no significant sensitivity to the dynamical state of the cluster. The chosen redshift bin size in the search is set to 0.06. The wider redshift binning compared to AMICO, by construction, regularises the contribution of galaxies with an excessively narrow PDZ, which would otherwise degrade the results.
@@ 3.2 Building the combined catalogue
The cluster detection algorithms were applied to each of the three Q1 fields. The clusters were selected in the redshift range , where the lower redshift bound corresponds to the nominal Euclid cluster catalogue requirement. We note that the upper redshift limit is lower than the one planned for subsequent data releases (), because photometry and photometric redshift estimation still have to be refined for these newly released data; the upper redshift limit for this sample is set by the poorly constrained behaviour of the Euclid photometric redshifts at , as shown in Figs. 2 and 3. In addition, as described in Sect. 2.2, the provisional requirement that all objects in the galaxy catalogues be detected in the VIS instrument necessarily restricts us to lower redshifts than what is anticipated for future data releases. While the principal catalogue is limited to , we nevertheless explore objects at in the joint catalogue in Sect. 4.7. This extension is undertaken with the knowledge of limitations in redshift accuracy, allowing us to assess the reliability of the cluster detections and highlight the most promising candidates in this regime.
@@
For this initial Q1 analysis, we ensured a consistent number density of detections between the two algorithms by adjusting the minimum S/N accordingly. Specifically, we used thresholds of S/N= () for AMICO and S/N= () for PZWav, to guarantee detections per square degree for each algorithm in the EDF-F and EDF-S (EDF-N). The different S/N thresholds applied to the southern and northern fields is due to the different photometric redshift quality, as discussed in Sect. 2.3. The absolute difference in values adopted for AMICO and PZWav is due to different definitions of the S/N by each finder. The S/N thresholds for each of the three fields were determined by visually inspecting a large number of detections to ensure an initial cluster sample that excludes clearly spurious detections. In the following analysis, we prioritise a joint sample of clusters in common between PZWav and AMICO to mitigate known issues in the input galaxy photometric redshifts. Focusing on the joint, high S/N detections for Q1 allows us to perform an initial internal validation of the cluster workflow, enabling us to identify any key issues that may impact the catalogue and subsequent cluster science applications.
To match the cluster detections between AMICO and PZWav, a physical distance of 1 Mpc was chosen, from which an angular search aperture was computed based on the mean cluster redshift of each pair. We imposed an additional constraint on any redshift difference of to minimise the number of erroneous matches to clusters in projection. If multiple matches were found, the highest S/N target was chosen. Fig. 4 shows the comparison of the S/N of the matched systems. The joint catalogue corresponds to the systems in the upper right part of the figure, which consists in 6.4 (EDF-N), 7.0 (EDF-S) and 7.9 (EDF-F) common detections out of the adopted nominal 15 detections per square degree for each algorithm. More of these detections are matched but were found at S/N lower than our thresholds (upper-left and lower-right quadrants of Fig. 4). The 35 highest S/N detections in the resulting joint catalogue are presented in Table 2. For each entry, we preserved the positions, redshifts, and S/N s derived by each detection algorithm, whereas the presented richness is computed at the position of the PZWav centres. Finally, the occurrence of fragmentation associated with the two detection algorithms is shown to be low over the overall mass range for clusters, with a more detailed assessment performed in Euclid Collaboration: Adam et al. (2019).
@@
The left panel of Fig. 7 compares the redshifts estimated by the two algorithms for the matched catalogue. The shaded pink region illustrates the matching interval, and we see that the redshift difference between matched objects is well within the interval, reinforcing the validity of the matching procedure. The right panel of Fig. 7 shows the internal consistency in the centring performance of the two algorithms.
@@
We computed the richnesses, , using only the photo- assignments for this study. We calculate the richness based on the PZWav cluster centre and redshift, but would obtain similar results by using the AMICO centre and redshift, based on the level of positional and redshift agreement shown in the two panels of Fig. 7. The combined richness and redshift distributions are presented in Fig. 9.
@@ 4.1 Searching for Euclid counterparts in external data
We cross-correlated the joint Q1 cluster catalogue with external meta-catalogues777A meta-catalogue is a combination of different source catalogues with cross-identification of clusters and homogenisation of key quantities, notably redshift and a mass proxy. and individual catalogues. We used meta-catalogues and catalogues primarily selected in X-ray – Meta-Catalogue of X-ray Clusters II (MCXC-II) described in Sadibekova et al. 2024, eROSITA (Bulbul et al. 2024; Kluge et al. 2024) – with the Sunyaev–Zeldovich (SZ) effect – Meta-Catalogue of SZ Clusters (MCSZ) described in Tarrio et al. (in preparation) – and in the optical – Dark Energy Survey (DES) Y1 RedMaPPer (Rykoff et al. 2016; Abbott et al. 2020), and the Abell catalogue (Abell 1958; Abell et al. 1989). We also used the joint X-ray and SZ selected catalogue ComPRASS (Tarr´ıo et al. 2019), the LC2 (Sereno 2015), and the Meta-Catalogue of Cluster Dispersions (MCCD) described in Euclid Collaboration: Melin et al. (in preparation) for clusters with high-quality velocity dispersion measurements.
@@
In total, we found 4/21/85 counterparts in the external catalogues for the 137/98/191 Euclid detections in the EDF-N, EDF-F, and EDF-S fields, respectively. In the EDF-F and the EDF-S, the matches are mainly from RedMaPPer and eROSITA. The EDF-N field is not covered by these two catalogues, which explains the significantly lower number of counterparts that are mainly from the MCXC-II. The total number of detections and matches from the external surveys that fall within the Q1 footprint are listed in footnote 9. A detailed breakdown of the individual matches to data sets per Q1 field can be found in Appendix B.